OpenStudy (anonymous):

A 80-kg man standing on a scale in an elevator notes that as the elevator rises, the scale reads 828 N. What is the acceleration of the elevator?

OpenStudy (anonymous):

F=mg+ (-mg) = may 828N=(the inside eqtn will cancel because it is the weight and the normal force and the man is not moving inside the elevator) = (80kg)ay 828N=(80kg)*ay 828N/80kg=ay 10.35 m/s^2

OpenStudy (anonymous):

Set up the equation like this: 1) ma = F in y-direction 2) Plug in given variables 3) (80kg)(?m/s^2) = 828 Newtons 4) Divide both side by 80 to get rid of that number and isolate acceleration. 5) Then you get 10.35 m/s^2 as your answer. (which is acceleration)